A risk-free investment of $6000 will return 3.5%. A risky $6000 investment has a 21% chance of defaulting and returning only $3500. What default premium must the risky investment offer?

A.10.5%

B. 12.0%

C. 11.6%     

D. 12.3%
Answer #1

Answer:

Given Return for risk free investment is 3.5%

So value of risk free investment =6000*(1+3.5%)=6210

Value of risky investment = Probability of defaulting * default return + (1-Probability of defaulting)*6000*(1+x)

Where X= return for risky investment

Value of risky investment=0.21*3500+0.79*6000*(1+x)=5475+4740X

So value of risk free investment=Value of risky investment

6210=5475+4740X

X=15.50%

So default risk premium = 15.50%-3.5%=12.00%